kpalm5989:

Which shows 60^2 − 40^2 being evaluated using the difference of squares method? ***I will give the first person a medal if they answer ASAP***

barstho:

What are the options?

kpalm5989:

A) 602 − 402 = (3600 + 1600)(3600 − 1600) = 10,400,000 B) 602 − 402 = 3600 − 1600 = 2,000 C) 602 − 402 = (60 − 40)2 = 202 = 400 D) 602 − 402 = (60 + 40)(60 − 40) = (100)(20) = 2,000

barstho:

Okay, so do you remember the difference of squares method?

kpalm5989:

Yes

barstho:

What is it?

kpalm5989:

DIFFERENCE OF PERFECT SQUARES METHOD: ^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^ Step 1: Factor out the GCF. Step 2: Determine whether the binomial is a difference of two squares. Step 3: Factor completely. ~ a2 − b2= (a + b)(a − b) Step 4: Check your factors.

barstho:

Perfect. So, it says: a^2-b^2 = (a+b)(a-b) We have, 60^2-40^2 (We know this is a difference of squares because there is a difference sign (-) and also they both are to the power 2) Now, use that formula on this.

kpalm5989:

Is the answer "D"... 60² − 40² = (60 + 40)(60 − 40) = (100)(20) = 2,000

barstho:

Yes!

kpalm5989:

Thank you so much!!!

barstho:

No problem (:

kpalm5989:

Now, how do I give you a medal?

barstho:

You already did lol

kpalm5989:

Thanks again!

barstho:

ofc